arccos x + 2arcsin sqrt3/2=pi/3
Answer by lwsshak3(11628)

arccos x + 2arcsin sqrt3/2=pi/3
arccos x + 2π/3=π/3
arccos x=-π/3
no solution, not in domain of arccos which is: (0,π)